# Interactive Identification of Granular Materials using Force Measurements

### v1

@@ IV-B Qualitative data analysis
To gain deeper insights into the nature of the granular materials and how they differ, we conducted a comprehensive qualitative data analysis. The analysis revealed several differences, supporting our hypothesis that the proposed interactive identification framework can effectively differentiate between these materials. Figure 2 illustrates typical F/T measurements obtained from interactions with each material. It is evident that there are observable differences not only in the magnitude of the measurements but also in their high-frequency content. Furthermore, we can observe differences in the signal dynamics, that is, in the way the signals evolve from the initial samples toward the end of the signal. Although it is not possible to comprehensively analyze all materials separately within a limited space, in the following, we will discuss the distinctive features at a general level while highlighting the most interesting examples.
@@
Figure 2: Materials and examples of their typical F/T measurements (zoom in for a better view)
@@ IV-B1 Signal magnitude
The most straightforward distinguishable factor between the different materials is the magnitude of the measurements, referred to as the signal magnitude. The signal magnitude indicates how much force is needed at each point in time during the interaction to maintain the constant end-effector velocity. Thus, the signal magnitude reflects information about the material’s flowing resistance, which correlates with its density and friction properties. Figure 2 highlights the differences between the materials. For instance, the average magnitude of high-density materials, like fine sugar, is significantly higher than that of lower-density materials, like wheat flour. However, relying solely on signal magnitude appears inadequate for identification, especially as the number of materials increases. This is evidenced, for example, by the similarities in the average signal magnitude of materials such as oat flakes and breadcrumbs.
@@ IV-B2 Frequency content
Figure 2 also highlights distinct frequency-related patterns exhibited by various materials. For example, rice yields low-amplitude, rapid zig-zag patterns, whereas wheat flour demonstrates smoother behavior. Fine sugar shows consistent and relatively smooth, high-amplitude wave patterns. Macaroni, on the other hand, produces measurements with irregular and sharp turns. As observed, materials with larger grain sizes or high friction, stemming from irregular grain shapes or surface properties, tend to produce sharper and more irregular patterns in the measurements. We hypothesize these patterns are associated with the grain size and shape, as well as the slide and rolling friction of the grains. These properties determine how smoothly the robot end-effector can traverse the material.
Although the patterns affect the signals’ frequency domain representations, pure frequency domain signals are noisy and lack distinguishable patterns. We will also show in the section V that classification results using frequency domain representations are poor.
@@ IV-B3 Signal dynamics
Signal dynamics refer to the qualitative patterns of how the signals develop from the beginning to the end. The signals, in general, exhibit non-stationarity, with the majority showing an increasing trend. This is expected, since the end-effector pushes the material forward during movement, causing the material to compress between the end-effector and the back edge of the container. Thus, a higher force is needed towards the end of the motion trajectory to maintain a constant velocity.
Some of the materials also display unique dynamics. For example, rice produces an easily identifiable high peak at the beginning of the signal, after which the signal descends close to zero and starts to gradually increase again towards the end. Another interesting, unique pattern can also be seen for clay granules. For clay granules, after an initial force increase, the readings take negative values. We associate this behavior with the dynamics of the material flow as the material moves away from the end-effector’s path and returns to fill the groove.
In summary, our analysis demonstrates that direct interaction with various granular materials yields meaningful force measurements. Distinct characteristics emerge for different materials in both the time and frequency domain. The provided analysis gives valuable insights for designing high-quality feature spaces that not only facilitate the classification of granular materials but also enable the estimation of more complex material properties, such as friction coefficients or flow dynamics, in the future. Based on our analysis, we hypothesize that a high-quality feature space for accurately distinguishing granular materials must consider the aforementioned characteristics of the measurements, including signal magnitude, frequency content, and signal dynamics.

### v2

@@ IV-B Qualitative data analysis
To gain deeper insights into the nature of the granular materials and how they differ, we conducted a comprehensive qualitative data analysis. The analysis aimed to examine whether the materials have enough differences to be classified and what the features distinguishing them are. The most significant forces were recorded along the x-axis and z-axis, while other forces remained closer to zero and less meaningful. Interestingly, forces along the z-axis tend to somewhat mirror those of the x-axis. Figure 2 illustrates typical F/T measurements obtained from interactions with each material. This figure presents the forces exclusively along the x- and z-axis to facilitate the inclusion of mean-variance plots while maintaining clarity and avoiding visual clutter. It can be observed that there are differences in both the magnitude of the measurements and their high-frequency content. Furthermore, we can observe differences in the signal dynamics, that is, in the way the signals evolve from the initial samples toward the end of the signal. Although it is impossible to comprehensively analyze all materials separately within a limited space, in the following, we will discuss the distinctive features at a general level while highlighting the most interesting examples.
@@
Figure 2: Materials and examples of their F/T measurements on the x- and z-axis, together with mean-variance plots (zoom in for a better view)
@@ IV-B1 Signal magnitude
The most straightforward distinguishable factor between the different materials is the magnitude of the measurements, referred to as the signal magnitude. The signal magnitude indicates how much force is needed at each point in time during the interaction to maintain the constant end-effector velocity. Thus, the signal magnitude reflects information about the material’s flowing resistance, which correlates with its density and friction properties. Figure 2 highlights the differences between the materials. For instance, the average magnitude of high-density materials, like fine sugar, is higher than that of lower-density materials, like sunflower seeds. However, relying solely on signal magnitude appears inadequate for identification, especially as the number of materials increases. This is evidenced, for example, by the similarities in the average signal magnitude of materials such as oat flakes and breadcrumbs. We evaluate the importance of the signal magnitude information empirically in Section V by comparing the classification results achieved with raw signals to the results with signals that are normalized between 0 and 1. We also compare the results when normalized and non-normalized signals are used together with the proposed HFMH feature space.
@@ IV-B2 Frequency content
Figure 2 also highlights distinct frequency-related patterns exhibited by various materials. For example, rice yields low-amplitude, rapid zig-zag patterns, whereas wheat flour demonstrates smoother behavior. Fine sugar shows consistent and relatively smooth, high-amplitude wave patterns. Macaroni, on the other hand, produces measurements with irregular and sharp turns. As observed, materials with larger grain sizes or high friction, stemming from irregular grain shapes or surface properties, tend to produce sharper and more irregular patterns in the measurements. These properties determine how smoothly the robot end-effector can traverse the material.
Although the patterns affect the signals’ frequency domain representations, pure frequency domain signals seem to lack distinguishable patterns. We will also show in the section V that classification results using pure frequency domain representations are sub-optimal. For this reason, our proposed method does not utilize the raw frequency information but instead uses a more compact presentation of the signal energy in high frequencies, as described in Section III.
@@ IV-B3 Signal dynamics
Signal dynamics refer to the qualitative patterns of how the signals develop from the beginning to the end. The signals, in general, exhibit non-stationarity, with the majority showing an increasing trend on the x-axis (decreasing on the z-axis after the initial peak). For forces along the x-axis, a rising trend is expected since the end-effector pushes the material forward during movement, causing the material to compress between the end-effector and the back edge of the container. Thus, a higher force is needed towards the end of the motion trajectory to maintain a constant velocity.
Some of the materials also display unique dynamics. For example, rice produces an easily identifiable high peak on the x-axis at the beginning of the signal, after which the signal descends close to zero and starts to gradually increase again towards the end. Another interesting pattern can be seen for clay granules and other clay granule-like materials (cat litter and, to some extent, potting gravel). For clay granule-like materials, after an initial force increase on the x-axis, the readings take negative values. We associate this behavior with the dynamics of the material flow as the material moves away from the end-effector’s path and returns to fill the groove.
In summary, our analysis demonstrates that direct interaction with various granular materials yields meaningful force measurements. Distinct characteristics emerge for different materials in both the time and frequency domain. The analysis in this section is also the basis for the design of the proposed feature space representation in this work. Furthermore, the provided analysis gives valuable insights for designing high-quality feature spaces that not only facilitate the classification of granular materials but also enable the estimation of more complex material properties, such as friction coefficients or flow dynamics, in the future.
@@ V Experiments and Results
Features Classification accuracy ($\pm$ SD)
Raw signal + HFMH 0.9705 ($\pm$ 0.0098)
Raw signal + DFT 0.7712 ($\pm$ 0.0426)
Norm. raw signal + HFMH 0.9303 ($\pm$ 0.0155)
Norm. raw signal + DFT 0.7686 ($\pm$ 0.0450)
HFMH 0.8705 ($\pm$ 0.0163)
Raw signal 0.7614 ($\pm$ 0.0265)
Raw hist. 0.7072 ($\pm$ 0.0434)
Norm. raw signal 0.8731 ($\pm$ 0.0251)
DFT 0.7621 ($\pm$ 0.0371)
PCA 0.7527 ($\pm$ 0.0373)
PCA + RICA 0.7602 ($\pm$ 0.0288)
TABLE I: Classification accuracy for different feature spaces. The higher, the better.